OpenStudy (anonymous):

A line has a slope of –5 and a y-intercept of (0, 3). What is the equation of the line that is perpendicular to the first line and passes through the point (3, 2)?

OpenStudy (texaschic101):

your first line has a slope of -5. You are looking for a perpendicular line, so you need the negative reciprocal of that slope. All that means is " flip " the slope and change the sign. So the negative reciprocal of -5, or -5/1, is 1/5 (see how I flipped the slope and changed the sign) now we will use y = mx + b slope(m) = 1/5 (3,2) x = 3 and y = 2 we have y and x and m, so we need to find b now lets sub 2 = 1/5(3) + b 2 = 3/5 + b 2 - 3/5 = b 10/5 - 3/5 = b 7/5 = b so your perpendicular line is : y = 1/5x + 7/5

OpenStudy (anonymous):

thank you!

OpenStudy (texaschic101):

sure thing :)